# Systematic study of confinement induced effects on atomic electronic structure

### v1

@@ 4.2.3 Comparison to literature radii
Numerically reliable ionization radii have been previously published by Boeyens3, Sen et al.13, and Garza et al.14. We compare our ionization radii from calculations based on the $\Delta$SCF method (section 4.2.1) and Janak’s theorem (section 4.2.2) with relaxed configurations against the earlier calculations in fig. 3.
@@
We first note that the values of Boeyens3 are systematically—and considerably—larger than the results of any of the other calculations for $Z\geq 6$. These surprisingly large differences are likely explained by the lack of electron correlation in Hartree–Fock theory, which is known to lead to overestimation of atoms’ size compared to high-level wave function calculations, with DFT yielding close agreement with accurate theoretical reference values.50 The closeness of DFT densities with multideterminantal densities has also been noted on previously by Ortiz-Henarejos and San-Fabián51, for example.
@@
Our calculations based on Janak’s theorem without the TSA systematically predict larger ionization radii than our $\Delta$SCF calculations, but they are still much smaller than the Hartree–Fock values of Boeyens3. Our $\Delta$SCF calculations are in excellent agreement with the limited data reported by Sen et al.13 in the spin-restricted formalism, who also found good agreement between their TSA method and $\Delta$SCF calculations in their work. Interestingly, our spin-polarized data appears to be in better agreement with the calculations of Sen et al.13 than our spin-restricted calculations. As was discussed in section 4.2.2, Sen et al.13 did not include correlation in their calculations, which likely explains the differences between our data.
@@
Our $\Delta$SCF calculations are also in good agreement for the 9 atoms studied by Garza et al.14 with (apparently $\Delta$SCF) calculations based on the spin-unrestricted local-density approximation.
@@
Since our $\Delta$SCF calculations yield ionization radii in good agreement with previous literature and ionization potentials of unconfined atoms in good agreement with experimental values (section 4.1), we are confident that our calculations offer a reliable basis for further studies of confined atoms with elaborate wave function methods. We further note that the dip in energy for the group 16 elements, which was found in the ionization energies in section 4.1, is also visible in the ionization radii in fig. 3.
(a) Spin-restricted densities
(b) Spin-polarized densities
@@
Figure 3: Comparison of ionization radii computed in this work via $\Delta$SCF (blue circles) or Janak’s theorem (magenta squares) against the Hartree–Fock values of Boeyens3 (red triangles) and DFT values of Sen et al.13 (forest green diamonds) and Garza et al.14 (navy crosses). The calculations of this work were carried out either with spin-restricted (fig. 3(a)) densities, or spin-polarized (fig. 3(b)) densities.
@@ 4.3 Observed changes in electronic structure
Finally, we examine the behavior of the low lying states of the H–Xe atoms as a function of the hard-wall confinement radius $r_{c}\in[a_{0},10a_{0}]$. We pay special consideration to ground state crossings, which signify states that may be especially relevant for chemical bonding.
Because the various configurations’ energies span several orders of magnitude under the studied range of confinement radii, our graphical representation of the results rests on comparing the energy of the confined atom with electron configuration $i$ to that of the ground state configuration of the unconfined atom
$\Delta E_{i}(r_{0})=E_{i}(r_{0})-E_{0}^{\mathrm{unconfined}},$ (11)
where $i$ is the state in question. The plots are included in the SI, as we restrict the discussion in the main text to the observed qualitative changes in atomic electronic structure under confinement.
In addition to energies, we also report estimated atomic radii for various electronic configurations of the unconfined atoms. First, the location of the density maximum of the outermost orbital is a long-established estimator for the size of covalent atoms52
where $n_{i}(r)$ is the $i$th orbital density. Second, since the electron density decays in the far valence region, a density threshold can be used to estimate the atom’s van der Waals size53
$n(r_{\mathrm{vdW}})=\epsilon.$ (13)
Bader et al.53 employed the threshold 0.002 electrons/bohr3, while Boyd54 showed that the criterion 0.001 electrons/bohr3 reproduces the same relative radii of atoms as those obtained with an even smaller value for the threshold. Rahm et al. also employed the threshold 0.001 electrons/bohr3 in their study of atomic and ionic radii55, 56 and we chose to adopt this value of the threshold, $\epsilon=10^{-3}$, as well.
We discovered in our analysis that $r_{\mathrm{vdW}}$ comes out surprisingly small for some diffuse states, such as the $1s^{1}3d^{1}$ configuration of He. The $r_{\mathrm{max}}$ estimator predicts this state to be considerably more extended than the $1s^{1}1p^{1}$ or $1s^{2}$ configurations ($r_{\mathrm{max}}=9.12a_{0}$, $3.68a_{0}$, and $0.56a_{0}$, respectively). Yet, $r_{\mathrm{vdW}}$ for the $1s^{1}3d^{1}$ configuration is smaller ($r_{\mathrm{vdW}}=2.08a_{0}$) than that for the $1s^{1}2p^{1}$ ($r_{\mathrm{vdW}}=4.10a_{0}$) or $1s^{2}$ ($r_{\mathrm{vdW}}=2.60a_{0}$) configurations.
To remedy this behavior, we introduce a third metric for atomic size by including the factor $r^{2}$ of the volume element in eq. 13 as
$r^{2}n(r_{\mathrm{r}^{2}\mathrm{vdW}})=\epsilon.$ (14)
This definition is somewhat more physical, since a spherical shell farther away from the nucleus can contain many more electrons, and thus having a small $n(r)$ does not alone guarantee that most of the electron density of the atom has been enclosed within $r_{\mathrm{r}^{2}\mathrm{vdW}}$. Due to the inclusion of the $r^{2}$ factor, eq. 14 appears to produce the largest estimates for atomic radii.
We note here that the radial density of electrons (the number of electrons found at $[r,r+{\rm d}r]$) is associated with an additional factor of $4\pi$. While the orbital density maximum in eq. 12 does not rely on a threshold and leads to the same radii whether or not this factor is included in the definition of $n(r)$, the threshold $\epsilon$ in eqs. 13 and 14 does depend on whether or not this factor has been included. Our implementation does not include the factor $4\pi$ in $n(r)$.
We provide a thorough analysis of the progression of the low lying states in hard-wall confinement in the SI. The analysis in the SI includes the three estimates of atomic radius for each configuration of the unconfined atom ($\epsilon=0.001$ for both $r_{\mathrm{vdW}}$ and $r_{\mathrm{r}^{2}\mathrm{vdW}}$), as well as their excitation energies relative to the ground state. We report the key observations from this analysis in the following.
@@
We do not observe ground state crossings for the H or He atoms in the studied range of confinement radii. The $1s$ orbital remains occupied over the whole domain of studied confinement radii. Interestingly, both atoms have configurations with an occupied $3d$ orbital which are stable with respect to ionization in the unconfined atom (the configuration has a lower energy than that of the cation’s ground state), even though the $3d$ orbital is extremely diffuse with $\langle r\rangle=\mathcal{O}(10a_{0})$.
@@
The valence electrons of the Li–O atoms occupy the $2s$ orbital in the unconfined atom. Contrary to the H and He atoms, we observe ground state crossings where the $2s$ electrons hop to the $2p$ orbital as the confinement radius is decreased; for the Li atom this happens already at $r_{c}=3a_{0}$, but the location of the crossover point $r_{c}$ decreases in increasing charge and occurs for the O atom at $r_{c}=a_{0}$. We also note that the Li and O atoms have configurations with an occupied $3d$ orbital that are stable with respect to ionization in the unconfined atom, although the $3d$ orbital is again extremely diffuse with $\langle r\rangle=\mathcal{O}(10a_{0})$, similarly to H and He.
@@
The valence electrons in the F atom occupy the $2s$ and $2p$ orbitals. One of the $2s$ electrons hops to the $2p$ orbital at $r_{c}=a_{0}$, which represents extreme confinement, thus fully filling the $2p$ orbital. Similarly to Li and O, F has a configuration with an occupied $3d$ orbital already in the unconfined atom, but the orbital is equally diffuse.
@@
The Ne atom does not exhibit ground state crossings in the studied range of confinement radii. We note that also Ne has a configuration with an occupied $3d$ orbital which is stable with respect to ionization in the unconfined atom.
@@
The valence electrons of Na and Mg occupy the $3s$ orbital, which becomes unfavourable in confinement, like the $2s$ orbital in Li and Be. While in Li and Be the electrons hopped from $2s$ to $2p$, the $3s$ electrons of Na and Mg hop to the $3d$ orbital around $r_{c}\approx 2a_{0}$, instead, even though the $3d$ orbital has a larger extent than the $3p$ orbital in the unconfined atom. This suggests that the $3d$ orbital can be relevant for chemical bonding.
We observe that the [Ne]$4f$ configuration of the Na atom is stable with respect to ionization in the unconfined atom, and flips below the initial ground state in extreme confinement (close to $r_{c}=a_{0}$). The $4f$ orbital is highly diffuse in the unconfined atom ($\langle r\rangle=\mathcal{O}(20a_{0})$). Also the state [Ne]$3d^{1}4f^{1}$ of Mg flips below the initial ground state close to $r_{c}=a_{0}$.
@@
The valence electrons of the unconfined Al–Ar atoms occupy the $3s$ and $3p$ orbitals. Unlike the atoms in the previous period, we see that the $np$ orbital becomes unfavourable in confinement, gradually transferring all electrons to the $3d$ orbital when the confinement radius is decreased, the transition occurring between $r_{c}=2a_{0}$ and $r_{c}=1.5a_{0}$ (strong confinement). As the confinement radius is further decreased, both the $3s$ electrons follow. The behavior is thus different from that observed above for Li–Ne, likely because the $3d$ orbital is too high to be accessible in the lighter atoms.
@@
Similarly to Na and Mg, the $ns$ valence electrons of the K and Ca atoms hop to the $3d$ orbital around $r_{c}\approx 4.5a_{0}$; the $3d$ orbital is thus already easily accessible, and likely relevant in chemistry. Further ground state crossings happen close to $r_{c}=a_{0}$ (extreme confinement), where the [Ar] core configuration is opened up and all the $3s$ and $3p$ electrons also move to the $3d$ orbital. We observe that both atoms have a configuration with an occupied $3d$ orbital that is bound relative to ionization in the gas-phase atom.
@@
The transition metals feature strong competition between the $ns$ and $(n-1)d$ orbitals, which can be observed from the small excitation energies in the unconfined atom. The states [Ar]$3d^{n}$, $n\leq 10$ are stable with respect to ionization already in the unconfined atom, which explains why ground state crossings happen already in weak confinement.
The $4s$ electrons shift to the $3d$ orbital around $r_{c}=4a_{0}$ in the Sc–Ni atoms, as was also observed for K and Ca. The Sc and Ti atoms exhibit further ground state crossings in extreme confinement (between $r_{c}=a_{0}$ and $r_{c}=2a_{0}$), where some $3s$ and $3p$ electrons also transfer to the $3d$ orbital.
@@
The unconfined Cu and Zn atoms already have a fully filled $3d$ orbital in their ground states, thus the $4s\rightarrow 3d$ shift observed earlier in $3d$ the transition metal series cannot happen for these atoms. The $4s$ electrons hop to the $4f$ orbital in extreme confinement (around $r_{c}=1.5a_{0}$), instead.
@@
The ground state configurations of the unconfined Ga–Kr atoms have filled $4s$ and $3d$ orbitals and a variably filled $4p$ orbital. The $4s$ and $4p$ electrons hop to the $4f$ orbital as the confinement radius is decreased (between $r_{c}=1.5a_{0}$ and $r_{c}=a_{0}$), leading to a [Ar]$3d^{10}4f^{n}$ ground state in extreme confinement.
@@
Like the alkali and earth alkali atoms in the previous period, the valence $ns$ electrons of the Rb and Sr atoms hop to the $(n-1)d$ orbital already around $r_{c}=5a_{0}$. However, unlike K and Ca, we also see a $4d\rightarrow 4f$ transfer when the confinement radius is further decreased to $r_{c}\approx 1.2a_{0}$, which represents extreme confinement. The $4s$ and $4p$ electrons also transfer to the $4f$ orbital below $r_{c}=1.2a_{0}$.
@@
The valence electrons of the unconfined Y–Rh atoms occupy the $ns$ and $(n-1)d$ orbitals in their ground states. The $5s$ valence orbital becomes unfavourable in confinement, and the electrons hop to the $4d$ orbital at confinement radii varying from $r_{c}=6.1a_{0}$ to $r_{c}=4.2a_{0}$, that is, already in weak confinement. Again, like for Cu–Sr, we observe electron transfers to the $4f$ orbital, and the $4d$ orbital is partially depopulated after the $5s$ orbital in confinement radii below $r_{c}=1.5a_{0}$, which already represents extreme confinement. Y and Zr transfer electrons from the $4p$ core orbital to the $4f$ orbital close to $r_{c}=a_{0}$.
@@
The Pd atom has a particularly stable gas-phase ground state ([Kr]$4d^{10}$). We observe no ground state crossings for Pd in the studied range of confinement radii.
@@
Ag and Cd already have a filled $4d$ orbital, while the $5s$ electrons hop to the $4f$ orbital around $r_{c}=2a_{0}$.
@@
The In–Xe atoms have filled $5s$ and $4d$ orbitals and a $5p$ orbital with variable filling in the unconfined atom. The $5p$ electrons start to gradually shift to the $4f$ orbital around $r_{c}=2a_{0}$, after which both the $5s$ electrons follow analogously to Ga–Kr.

### v2

@@ 4.2.3 Comparison to literature radii
Numerically reliable ionization radii have been previously published by Boeyens5, Sen et al.15, and Garza et al.16. We compare our ionization radii from calculations based on the $\Delta$SCF method (section 4.2.1) and Janak’s theorem (section 4.2.2) with relaxed configurations and the PBE functional against the earlier calculations in fig. 7. Analogous plots for the PW92 and r2SCAN functionals can be found in the SI. We note that the dip in energy for the group 16 elements, which was found in the ionization energies in section 4.1, is also visible in the ionization radii in fig. 7.
@@
We first note that the analysis of Boeyens5 was based on non-relativistic Hartree–Fock calculations for $1\leq Z\leq 102$, and $r_{+}$ was identified via Koopmans’ theorem45 by finding the zero of the orbital energy of the outermost valence electron shell. However, as discussed in section 1, Boeyens5 did not employ a proper hard-wall potential, and did not consider the changing nature of the ground state configuration as a function of confinement. The values of Boeyens5 are systematically and considerably larger than the results of any of the other calculations for $Z\geq 15$. These surprisingly large differences are likely explained by the lack of electron correlation in Hartree–Fock theory, which is known to lead to overestimation of atoms’ size compared to high-level wave function calculations, with DFT yielding close agreement with accurate theoretical reference values.66 The closeness of DFT densities with multideterminantal densities has also been noted on previously by Ortiz-Henarejos and San-Fabián67, for example.
@@
Our calculations based on Janak’s theorem without the TSA systematically predict larger ionization radii than our $\Delta$SCF calculations, but they are still much smaller than the Hartree–Fock values of Boeyens5. Our $\Delta$SCF calculations are in excellent agreement with the limited data reported by Sen et al.15 in the spin-restricted formalism, who also found good agreement between their TSA method and $\Delta$SCF calculations in their work. Interestingly, our spin-polarized data appears to be in better agreement with the calculations of Sen et al.15 than our spin-restricted calculations, as the additional comparisons included in the SI show, even though the calculations of Sen et al.15 were spin-restricted. As was discussed in section 4.2.2, Sen et al.15 did not include correlation in their calculations, which likely explains the differences between our data.
@@
Our $\Delta$SCF calculations are also in good agreement for the 9 atoms studied by Garza et al.16 with spin-polarized $\Delta$SCF calculations with the PW92 functional.
@@
Since our $\Delta$SCF calculations were found to produce ionization potentials of unconfined atoms in good agreement with experimental values (section 4.1), and also yield ionization radii in good agreement between dissimilar density functional approximations as well as with previous literature, we are confident that our calculations offer thus a reliable basis for studies of confined atoms, and a reliable ground for further studies with elaborate wave function methods.
@@
Figure 7: Comparison of ionization radii computed in this work with spin-polarized densities and the PBE density functional via $\Delta$SCF (blue circles) or Janak’s theorem (green squares) against the Hartree–Fock values of Boeyens5 (red triangles) and DFT values of Sen et al.15 (magenta diamonds) and Garza et al.16 (orange crosses).
4.3 Estimates for atomic radii
Before proceeding to the in-depth analysis of the changes of atoms’ electronic structure in confinement, we return to the question of the atomic radii discussed above in section 2.4. Because the present study involves many configurations for each atom, many not being bound with respect to ionization in the unconfined atom, it is useful to be able to estimate the resulting atoms’ size. While the covalent size estimate of Slater50 in eq. 9 is parameter-free, the two vdW radius estimates of eqs. 10 and 11 feature a threshold parameter.
Bader et al.51 employed the threshold 0.002 electrons/bohr3 in eq. 10, while Boyd68 showed that the criterion 0.001 electrons/bohr3 reproduces the same relative radii of atoms as those obtained with an even smaller value for the threshold. Rahm et al. also employed the threshold 0.001 electrons/bohr3 in their study of atomic and ionic radii69, 70 and we chose to adopt this value of the threshold, $\epsilon=10^{-3}$ in eq. 10, as well. We note that Smith et al.71 have also used eq. 10 recently to estimate vdW radii, opting to use a threshold of 0.0015 electrons/bohr3, instead.
We discovered an issue in our analysis with the $r_{\rho}$ estimate of eq. 10 for the vdW radius for the $3d^{1}$ state of the H atom. As only $s$-type orbitals can have a finite value at the nucleus,53 and as the $3d$ orbital in question turns out to be extremely diffuse in character, the density criterion eq. 10 is not satisfied with any radius. As a result, $r_{\rho}$ is not defined for the employed value of $\epsilon$.
In addition, $r_{\rho}$ for the $1s^{1}3d^{1}$ configuration of He turns out to be smaller ($r_{\rho}=2.10a_{0}$ for spin-unrestricted PBE) than that for the $1s^{1}2p^{1}$ ($r_{\rho}=4.05a_{0}$) or $1s^{2}$ ($r_{\rho}=2.68a_{0}$) configurations, even though basic physical insight would say that the $1s^{2}$ ground state configuration has to be more compact than those where an electron is excited from the $1s$ shell to the $2p$ shell, not to mention to the $3d$ shell.
We also find that the criterion of eq. 10 is overly sensitive to the numerical discretization: a finite element calculation for the exact ground state of the unconfined hydrogen atom reveals a significant sensitivity to numerical noise, as shown in fig. 8. We find it highly likely that similar issues are also present in the Gaussian-basis calculations of refs. 69, 70, 71 that relied on the estimate of eq. 10: small changes to the Gaussian basis likely result in considerable changes in the estimated radius, since the calculations of fig. 8 which all reproduce the exact ground state energy $-1/2E_{h}$ to sub-$\mu E_{h}$ accuracy exhibit considerable differences in the estimate of the vdW radius.
Note that the ANO-RCC basis set employed in refs. 69 and 70 produces an analogous Hartree–Fock energy of $-0.49998374E_{h}$ (basis set obtained from the Basis Set Exchange72 and calculation performed with ERKALE73), and thus exhibits a much larger truncation error of $16.3\mu E_{h}$ for the unconfined hydrogen atom. Smith et al.71 examined various Gaussian basis sets and variations in the order of $0.01a_{0}$ can be observed in their data, as well.
[…]
The data in fig. 8 show that the proposed $r_{\epsilon}$ estimate of eq. 11 is numerically well-behaved, yielding excellent agreement with the radius determined from the analytical solution, thus successfully addressing the issues with undefined vdW radii and numerical stability in the $r_{\rho}$ estimate of eq. 10. This criterion also easily resolves the issue with the configurations of He, agreeing with physical insight that the $1s^{2}$ ground state ($r_{\epsilon}=2.26a_{0}$ for spin-unrestricted PBE) is more compact than the $1s^{1}2p^{1}$ ($r_{\epsilon}=8.07a_{0}$) or $1s^{1}3d^{1}$ ($r_{\epsilon}=15.48a_{0}$) excited states, the latter of which also show a significant difference in diffuse character.
The $r_{\epsilon}$ estimates for the states of He are also in excellent qualitative agreement with the covalent radius estimates $r_{\rm max}$ of eq. 9: both $r_{\epsilon}$ and $r_{\rm max}$ predict the size order $1s^{2}$ < $2p^{2}$ < $1s^{1}2p^{1}$ < $1s^{1}3d^{1}$, while the $r_{\rho}$ estimate of eq. 10 predicts a thoroughly dissimilar order $1s^{1}3d^{1}$ < $1s^{2}$ < $1s^{1}2p^{1}$ < $2p^{2}$.
We note again here that the $1s^{1}3d^{1}$ excited state is so diffuse in the unconfined He atom that it is likely not converged to the complete basis set limit in our calculations, as we chose not to converge the value of the practical infinity $r_{\infty}$ given that excited states of the unconfined atom are not the focus of this work.
Figure 8: The dependence of the radial estimates $r_{\rho}$ and $r_{\epsilon}$ defined by eqs. 10 and 11, respectively, on the number of radial elements $N_{\text{elem}}$ used in the discretization. The calculations are performed for the ground state of the hydrogen atom, using unrestricted Hartree–Fock theory, and the error in the total energy compared to the analytical value $-0.5E_{h}$ is smaller than $10^{-7}E_{h}$ in all calculations shown in the figure. The estimates from eq. 11 agree with the analytical value $r_{\epsilon}\approx 2.8815126965663684390a_{0}$ to a precision better than one part per million.
@@ 4.4 Observed changes in electronic structure
In the following, we will discuss the key observations from the systematic analysis of all the atoms H–Xe. We focus the discussion in the main text on the PBE functional due to its insurmountable popularity in the literature, but note that the behaviors of various atoms are qualitatively independent of the functional, as was already demonstrated for the Ca atom in section 4.2. Thorough analyses for all functionals are available in the SI.
In addition to energies of the configurations as a function of confinement, we also report the estimated atomic radii for the configurations of the unconfined atoms with the methodology discussed in sections 2.4 and 4.3 in the SI. We note again the limitations of the previously used $r_{\rho}$ estimate of eq. 10 that was discussed above in section 4.3, in that the reported $r_{\rho}$ values exhibit numerical noise. Furthermore, the excitation energies of each configuration relative to the ground state are also included in the SI for the case of the unconfined atom.
We note here that we were unable to converge the calculations of some configurations with the r2SCAN functional. We tentatively attribute this to the numerical ill-behavedness of the functional74 and to the slow convergence of meta-GGA functionals in general with respect to the number of radial elements.38 We have excluded all non-converged calculations from our analysis.
The in-depth analysis of the data included in the SI led us to the following observations on the behavior of the H–Xe atoms, which can be grouped into the following classification based on the atom’s position in the periodic table.
@@
Similarly to previous studies, we do not observe ground state crossings for the H or He atoms in the studied range of confinement radii. The $1s$ orbital remains occupied over the whole domain of studied confinement radii. Interestingly, both atoms have configurations with an occupied $3d$ orbital which are stable with respect to ionization in the unconfined atom (the configuration has a lower energy than that of the cation’s ground state), even though the $3d$ orbital is extremely diffuse as $r_{\epsilon}\approx 15a_{0}$ for the corresponding configuration.
@@
The valence electrons of the Li–O atoms occupy the $2s$ orbital in the unconfined atom. Contrary to the H and He atoms, we observe ground state crossings where the $2s$ electrons hop to the $2p$ orbital as the confinement radius is decreased; for the Li atom this happens already at $r_{c}=3a_{0}$, but the location of the crossover point $r_{c}$ decreases in increasing charge and occurs for the O atom at $r_{c}=a_{0}$.
The electron shift for Li was also predicted by Rahm et al.22, while the $2s^{2}\to 2p^{2}$ transition of the C atom was reported by Pašteka et al.19. As far as we are aware, the analogous transitions for the Be, B, N, and O atoms have not been discussed in previous literature.
We further note that the Li and O atoms have configurations with an occupied $3d$ orbital that are stable with respect to ionization in the unconfined atom, although the $3d$ orbital is again extremely diffuse as $r_{\epsilon}\approx 15a_{0}$, as in the case of H and He.
@@
The valence electrons in the F atom occupy the $2s$ and $2p$ orbitals. One of the $2s$ electrons hops to the $2p$ orbital at $r_{c}=a_{0}$, which represents extreme confinement, thus fully filling the $2p$ orbital. This electron transition does not appear to have been previously reported in the literature. Similarly to Li and O, F has a configuration with an occupied $3d$ orbital already in the unconfined atom, but the orbital is equally diffuse.
@@
Confirming previous studies, the Ne atom does not exhibit ground state crossings in the studied range of confinement radii. We note that also Ne has a configuration with an occupied $3d$ orbital which is stable with respect to ionization in the unconfined atom.
@@
The valence electrons of Na and Mg occupy the $3s$ orbital, which becomes unfavourable in confinement, like the $2s$ orbital in Li and Be. While in Li and Be the electrons hopped from $2s$ to $2p$, the $3s$ electrons of Na and Mg hop to the $3d$ orbital around $r_{c}\approx 2a_{0}$, instead, even though the $3d$ orbital has a larger extent than the $3p$ orbital in the unconfined atom, and the $3d$ orbital is not a valence orbital even in the Mg atom without confinement.75 These electron transitions, which occur at relatively strong confinement, have not been previously reported in the literature to our knowledge.
We observe that the [Ne]$4f$ configuration of the Na atom is stable with respect to ionization in the unconfined atom, and flips below the initial ground state in extreme confinement (close to $r_{c}=a_{0}$). The $4f$ orbital is highly diffuse in the unconfined atom ($r_{\epsilon}\approx 24a_{0}$). Also the state [Ne]$3d^{1}4f^{1}$ of Mg flips below the initial ground state close to $r_{c}=a_{0}$.
@@
The valence electrons of the unconfined Al–Ar atoms occupy the $3s$ and $3p$ orbitals. No electron shifts have been reported in previous studies for these atoms. However, we see that the $3p$ orbital becomes unfavourable in confinement, gradually transferring all electrons to the $3d$ orbital when the confinement radius is decreased, the transition occurring between $r_{c}=2a_{0}$ and $r_{c}=1.5a_{0}$ (strong confinement). As the confinement radius is further decreased, both the $3s$ electrons follow. The behavior is thus different from that observed above for Li–Ne, likely because the $3d$ orbital is too high to be accessible in the lighter atoms.
@@
Confirming the results of previous studies,22, 9, 17 analogously to Na and Mg, the $4s$ valence electrons of the K and Ca atoms hop to the $3d$ orbital around $r_{c}\approx 4.5a_{0}$. The $3d$ orbital is thus already easily accessible, and its relevance in chemistry has been well established in the literature:76, 77, 75, 78, 79, 80, 81 Ca is sometimes referred to in the literature as an “honorary transition metal”81 due to its covalent binding like a transition metal,77, 75, 78, 80 following similar work performed earlier on the heavier analogoues Cs and Ba which are not considered in this work.82
Further novel ground state crossings happen close to $r_{c}=a_{0}$ (extreme confinement), where the [Ar] core configuration is opened up and all the $3s$ and $3p$ electrons also move to the $3d$ orbital. We observe that both atoms have a configuration with an occupied $3d$ orbital that is bound relative to ionization in the unconfined atom.
@@
The $3d$ transition metals feature strong competition between the $4s$ and $3d$ orbitals, which can be observed from the small excitation energies in the unconfined atom. The states [Ar]$3d^{n}$, $n\leq 10$ are stable with respect to ionization already in the unconfined atom, which explains why ground state crossings happen already in weak confinement.
Confirming previous findings,22, 9 the $4s$ electrons shift to the $3d$ orbital around $r_{c}=4a_{0}$ in the Sc–Ni atoms, similarly to the observations made above for K and Ca. However, here we observe that the Sc and Ti atoms exhibit additional ground state crossings in extreme confinement (between $r_{c}=a_{0}$ and $r_{c}=2a_{0}$), where some $3s$ and $3p$ electrons also shift to the $3d$ orbital.
@@
The unconfined Cu and Zn atoms already have a fully filled $3d$ orbital in their ground states, thus the $4s\rightarrow 3d$ shift observed earlier in the $3d$ transition metal series cannot happen for these atoms. The $4s$ electrons hop to the $4f$ orbital in extreme confinement (around $r_{c}=1.5a_{0}$), instead, which does not appear to have been previously reported in the literature.
@@
The ground state configurations of the unconfined Ga–Kr atoms have filled $4s$ and $3d$ orbitals and a variably filled $4p$ orbital. The $4s$ and $4p$ electrons hop to the $4f$ orbital as the confinement radius is decreased (between $r_{c}=1.5a_{0}$ and $r_{c}=a_{0}$), leading to a [Ar]$3d^{10}4f^{n}$ ground state in extreme confinement. The favouring of the $4f$ orbital over the $4s$ and $4p$ orbitals in the confined Kr atom has been previosly reported by Garza et al.16.
@@
Like the alkali and earth alkali atoms in the previous period, the valence $5s$ electrons of the Rb and Sr atoms hop to the $4d$ orbital already around $r_{c}=5a_{0}$ confirming the previous observations of Guerra et al.17 and Rahm et al.22. However, unlike the lighter analogues K and Ca, we also see a $4d\rightarrow 4f$ shift when the confinement radius is further decreased to $r_{c}\approx 1.2a_{0}$, which represents extreme confinement. The $4s$ and $4p$ electrons also shift to the $4f$ orbital below $r_{c}=1.2a_{0}$, which, to the best of our knowledge, has not been reported in the literature so far.
@@
The valence electrons of the unconfined Y–Rh atoms occupy the $5s$ and $4d$ orbitals in their ground states. The $5s$ valence orbital becomes unfavourable in confinement, and the electrons hop to the $4d$ orbital at confinement radii varying from $r_{c}=6.1a_{0}$ to $r_{c}=4.2a_{0}$, that is, already in weak confinement, in agreement with previous literature.22, 9 Again, like for Cu–Sr, we observe further electron shifts to the $4f$ orbital, and the $4d$ orbital is partially depopulated after the $5s$ orbital in confinement radii below $r_{c}=1.5a_{0}$, which already represents extreme confinement. Y and Zr shift electrons from the $4p$ core orbital to the $4f$ orbital close to $r_{c}=a_{0}$. These additional electron shifts do not appear to have been reported in previous studies.
@@
It is well known that the unconfined Pd atom has a particularly stable ground state ([Kr]$4d^{10}$).22, 9 We observe no ground state crossings for Pd in the studied range of confinement radii.
@@
Ag and Cd already have a filled $4d$ orbital, while we observe previously unexplored electron shifts where $5s$ electrons hop to the $4f$ orbital around $r_{c}=2a_{0}$.
@@
The In–Xe atoms have filled $5s$ and $4d$ orbitals and a $5p$ orbital with variable filling in the unconfined atom. We observe that the $5p$ electrons start to gradually shift to the $4f$ orbital around $r_{c}=2a_{0}$, after which both the $5s$ electrons follow analogously to Ga–Kr, which does not appear to have been reported in the literature before now.